OpenStudy (anonymous):

2t/5=7/9 solve and report in lowest terms

OpenStudy (anonymous):

2t-y+3=19

OpenStudy (anonymous):

2t-y+3=9

OpenStudy (anonymous):

o nvr mind he got it

OpenStudy (anonymous):

the answer is 7/2 I just dont know how they got that

OpenStudy (anonymous):

cross multiply.

OpenStudy (anonymous):

9(2t)=5*7

OpenStudy (anonymous):

7/2 or 7/5?

OpenStudy (anonymous):

i think i copied the question wrong... its really 2t/5=7/5

OpenStudy (anonymous):

ohh lol

OpenStudy (anonymous):

2t(5)=5*7

OpenStudy (anonymous):

can you solve that?